OpenStudy (anonymous):

Given that a, b, and c are non-zero real numbers and a + b ≠ 0, solve for x. ax + bx - c = 0

OpenStudy (aum):

ax + bx - c = 0 factor x out of the first two terms: x(a + b) - c = 0 add c x(a + b) = c divide by (a + b) x = c / (a + b)

OpenStudy (anonymous):

thank you so much